# Inertial Torsion Noise in Matter-Wave Interferometers for Gravity Experiments

### v1

@@ IV Power Spectrum Density of Inertial Torsion Noise
There is also a peak of the PSD of ITN, of which the resonance frequency translates from $\Omega_{\rm rot}$ to $2\Omega_{\rm rot}$, which is a common property for a squared noise  [^9] This property can be understood as follows. Since $S_{\dot{\Theta}\dot{\Theta}}$ has a peak at $\Omega_{\rm rot}$, then it can be written as a $\delta$-function plus some small function $o(\omega)$: $S_{\dot{\Theta}\dot{\Theta}}(\omega)=A\delta(\omega-\Omega_{\rm rot})+o(\omega),$ where $A$ is the amplitude of the peak. Then the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}$ gives $\begin{split}S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)&=A^{2}\int\delta(\omega^{\prime}-\Omega_{\rm rot})\delta(\omega-\omega^{\prime}-\Omega_{\rm rot})\mathrm{d}\omega^{\prime}+o^{\prime}(\omega)\\ &=A^{2}\delta(\omega-2\Omega_{\rm rot})+o^{\prime}(\omega),\end{split}$ which means the peak position locates at $2\Omega_{\rm rot}$. In Appendix B, we offer another method to understand this property in the time domain.. In a small damping limit $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, the peak value of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is
$\begin{split}S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(2\Omega_{\rm rot})&=\frac{4\pi\gamma_{\rm rot}(k_{B}T)^{2}}{I^{2}}\frac{32\Omega_{\rm rot}^{4}+16\gamma_{\rm rot}^{2}\Omega_{\rm rot}^{2}}{16\gamma_{\rm rot}^{2}\Omega_{\rm rot}^{2}(4\Omega_{\rm rot}^{2}+\gamma_{\rm rot}^{2})}\\ &\approx\frac{2\pi(k_{B}T)^{2}}{\gamma_{\rm rot}I^{2}}.\end{split}$ (24)
The quality factor (Q-factor) of ITN can be computed according to the frequency-to-bandwidth ratio known as the full width at half maximum (FWHM). In particular, under the small-damping limit $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, i.e. the high-Q case, one can compute that  [^10] In particular, one needs to solve the equation $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)=S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(2\Omega_{\rm rot})/2$ to get the FWHM frequency $\omega_{1}$ and $\omega_{2}$, then the Q-factor is defined as $Q\equiv 2\Omega_{\rm rot}/\Delta\omega$ with $\Delta\omega\equiv\omega_{2}-\omega_{1}$. For ease of calculation, we only focus on the small-damping limit $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, corresponding to the high-Q case. According to the PSD of ITN (22) and the approximated formula (24) of the peak value, the equation for the FWHM frequency is $\frac{4\omega^{4}+4(\gamma_{\rm rot}^{2}-3\Omega_{\rm rot}^{2})\omega^{2}+16\Omega_{\rm rot}^{4}}{(\omega^{2}+\gamma_{\rm rot}^{2})(4\gamma_{\rm rot}^{2}\omega^{2}+(\omega^{2}-4\Omega_{\rm rot}^{2})^{2})}=\frac{1}{2}\cdot\frac{1}{2\gamma_{\rm rot}^{2}}.$
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(2\Omega_{\rm rot}\pm\gamma_{\rm rot})\approx\frac{\pi(k_{B}T)^{2}}{\gamma_{\rm rot}I^{2}}.$ (25)
It suggests that the FWHM frequency is $\omega=2\Omega_{\rm rot}\pm\gamma_{\rm rot}$. Thus, the frequency bandwidth is $\Delta\omega=2\gamma_{\rm rot}$ and the Q-factor for ITN is
$Q_{\rm ITN}=\frac{2\Omega_{\rm rot}}{2\gamma_{\rm rot}}=\frac{\Omega_{\rm rot}}{\gamma_{\rm rot}},$ (26)
which is the same as the Q-factor of $S_{\Theta\Theta}(\omega)$.
@@ V Dephasing Factor and Experiment Parameters Constrain
Figure 5: Dephasing factor $\Gamma$ as a function of the intrinsic torsion frequency $\Omega_{\rm rot}$. The parameters are chosen as $m_{0}=10^{-15}$ kg, $\eta=10^{4}$ T/m, $t_{a}=0.25$ s and $t_{e}=0$ s for the interferometer, and $\gamma_{\rm rot}=10^{-10}$ Hz, $I=10\,\rm kg\cdot m^{2}$ and $T=300$ K for the PSD of ITN. There is a resonance between the ITN and the transfer function with the first dominant peak at $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}\sim 2\pi\times 1$ Hz, and additional smaller peaks visible in the range upt to 10. In low frequency limit and high frequency limit, the dephasing factor $\Gamma$ is approximately independent with the intrinsic frequency $\Omega_{\rm rot}$.
Figure 6: Dephasing factor $\Gamma$ as a function of the damping factor $\gamma_{\rm rot}$. The transfer function is given in Eq. (13) with parameters $t_{a}=0.25$ s and $t_{e}=0$ s. Three different $\Omega_{\rm rot}$ are chosen in the low rotation frequency limit (blue line), the resonance frequency condition $\Omega_{\rm rot}\sim 2\pi/T=2\pi$ Hz (orange line), and the high rotation frequency limit (green line). The dephasing $\Gamma$ is proportional to $\gamma_{\rm rot}$ when $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, while it starts decreasing with respect to $\gamma_{\rm rot}$ when $\gamma_{\rm rot}$ becomes comparable or larger than $\Omega_{\rm rot}$.
@@
Based on the result of the transfer function in Eq. (13) or Eq. (14), and the power spectrum density (22) of the inertial torsion noise, the dephasing factor $\Gamma$ can be computed through the integral in Eq. (10). As this integral is non-trivial, we employ numerical methods to compute it. It is noteworthy that a resolution of $\omega_{\rm min}=2\pi/T_{\rm tot}$ exists as a cutoff in frequency space for the numerical calculation  [^11] This resolution arises from the periodic condition assumption of discrete Fourier transform (DFT). Suppose a signal $x(t)$ with a total sample time $T$ satisfies a periodic condition $x(t+T)=x(t),\ \forall t$. In frequency space, this condition can be written as $\int X(\omega)\mathrm{e}^{\mathrm{i}\omega(t+T)}\mathrm{d}(t+T)=\int X(\omega)\mathrm{e}^{\mathrm{i}\omega t}\mathrm{d}t.$ Since this equation holds for any $t$, we have the constraint that $\mathrm{e}^{\mathrm{i}\omega T}=1$, which implies that $\omega T=2k\pi$ for some integer $k$. Thus, a minimum frequency $\omega_{\rm min}=2\pi/T$ as a cut-off in frequency space, known as the frequency resolution.. Physically, this cutoff indicates that a low-frequency signal or noise will not be measurable by the experiment if the total experiment time is shorter than a single period of such signal or noise.
@@
Fig. 5 shows the dependence of $\Gamma$ on the intrinsic torsion frequency $\Omega_{\rm rot}$ of the experiment apparatus, where the parameters of $F(\omega)$ are chosen as $m_{0}=10^{-15}$ kg, $t_{a}=0.25$ s and $t_{e}=0$ s. As is shown, there is a resonance between the ITN and the transfer function near $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}=2\pi\times 1$ Hz and some harmonic resonances. Note that near the resonance peak, the precision of the numerical calculation is limited because the value is highly related to the sample points in frequency space.
@@
When $\Omega_{\rm rot}$ is much smaller than $2\pi\times 10^{-1}$ Hz, the peak will be below the frequency cut-off $\omega_{\rm min}$, so the dephasing factor will be tiny. For an actual experiment, the low intrinsic frequency limit is expected, so the parameters $d$, $l$, $\mathbb{G}$ and $I$ in (17) have to be carefully designed to ensure $\Omega_{\rm rot}$ much smaller than the resonance frequency $2\pi/T$ of the interferometer.
When $\Omega_{\rm rot}$ is much larger than $2\pi\times 10^{0}$ Hz, $\Gamma$ becomes approximately independent on $\Omega_{\rm rot}$. Fig. 6 shows the dependence of $\Gamma$ on the damping factor $\gamma_{\rm rot}$ caused by the thermal gas molecules. Three curves with different $\Omega_{\rm rot}$ correspond to the low rotation frequency limit, the high rotation frequency limit and the resonance frequency condition $\Omega_{\rm rot}\sim 2\pi/T$. As is shown in the figure, $\Gamma$ is proportional to $\gamma_{\rm rot}$ when $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, which can be read from Eq. (22). When $\gamma_{\rm rot}$ grows comparable to $\Omega_{\rm rot}$, $\Gamma$ decreases with respect to $\gamma_{\rm rot}$ increases.
Fig. 6 show the dependence of $\Gamma$ on the damping rate $\gamma_{\rm rot}$ of the gas collision. As is shown, when the damping rate is very low, i.e. $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, the dephasing factor $\Gamma$ is approximately proportional to $\gamma_{\rm rot}$ because the $\bar{S}$ in Eq. (22) is proportional to $\gamma_{\rm rot}$.
However, when $\gamma_{\rm rot}$ increases after a critical point $\gamma_{\rm rot}\sim\Omega_{\rm rot}$, the dephasing factor $\Gamma$ is approximately inverse to $\gamma_{\rm rot}$. This is because the PSD of ITN at the high damping rate limit $\gamma_{\rm rot}\gg\Omega_{\rm rot}$ is approximately
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)\approx\bar{S}\frac{4\omega^{4}+4\gamma_{\rm rot}^{2}\omega^{2}}{(\omega^{2}+\gamma_{\rm rot}^{2})(\omega^{4}+4\gamma_{\rm rot}^{2}\omega^{2})}\sim\frac{\bar{S}}{\gamma_{\rm rot}^{2}}.$ (27)
Since $\bar{S}\propto\gamma_{\rm rot}$, the PSD of ITN is proportional to $\gamma_{\rm rot}^{-1}$, so the dephasing factor $\Gamma$ decreases with respect to $\gamma_{\rm rot}$. Physically it can be interpreted as an overdamped oscillator. In particular, the apparatus’s dynamical equation in Eq. (18) describes a damped oscillator under a randomly driven force. When the damping rate is larger than a critical value $\gamma_{\rm rot}>\Omega_{\rm rot}/2$, the system decays with no oscillation, known as overdamped. In the overdamped region, as the damping rate increases, the system decays to the equilibrium faster, so the random force term affects the system less. Finally, the dephasing factor $\Gamma$ decreases as the damping rate $\gamma_{\rm rot}$ increases.
A final remark on Figs. 5 and 6 we note that the dephasing factor $\Gamma$ caused by the torsion noise from gas molecules collision is negligibly tiny. Two main reasons cause this. First, the thermal motion of gas molecules is proportional to a small thermal factor $k_{B}T\sim 10^{-21}$ J. Besides, ITN is a second-order effect, so every little factor in the PSD $S_{\Theta\Theta}(\omega)$ of the experiment apparatus will be squared for the dephasing factor $\Gamma$. Combining both effects, the dephasing factor $\Gamma$ is suppressed by a highly tiny factor $(k_{B}T)^{2}\sim 10^{-42}$, such that it does not exceed $10^{-30}$ in these two plots.
VI a generic source of torsion noise
In an actual experiment, there can be a number of sources of torsion noise. To model generic noises phenomenologically we can assume the Langevin equation of the form
$\ddot{\Theta}=-\Omega_{\rm rot}^{2}\Theta-\gamma\dot{\Theta}+\sqrt{A}\Theta_{\rm in},$ (28)
where the dissipation rate $\gamma$ and the amplitude of the random Gaussian force $\sqrt{A}$ is determined by the experiment (and we recall that $\Omega_{\rm rot}$ is the intrinsic frequency). Te PSD of the motion of the experiment box and the ITN can be computed from the dynamical equation following analogous steps as in the previous sections:
$S_{\Theta\Theta}(\omega)$ $=\frac{A}{(\Omega_{\rm rot}^{2}-\omega^{2})^{2}+\gamma^{2}\omega^{2}},$ (29)
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ $=A^{2}\frac{4\omega^{4}+4(\gamma^{2}-3\Omega_{\rm rot}^{2})\omega^{2}+16\Omega_{\rm rot}^{4}}{\gamma(\omega^{2}+\gamma^{2})(4\gamma^{2}\omega^{2}+(\omega^{2}-4\Omega_{\rm rot}^{2})^{2})}.$
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ has an additional $\gamma^{-2}$ dependence in comparison to the thermal case in Eq. (22). This difference arises because the amplitude $A=2\gamma_{\rm rot}k_{B}T/I$ of the external force for the thermal noise in Eq. (18) is proportional to the damping rate $\gamma$, which does not hold for the general case considered in this section. In this case, the dependence of $\Gamma$ on the damping rate $\gamma$ is shown as Fig. 7, where the amplitude $A$ is chosen as $10^{-10}\,\rm Hz^{3}$. As is shown, $\Gamma$ is proportional to $\gamma^{-1}$ and $\gamma^{-3}$ in the underdamped region $\gamma\ll\Omega_{\rm rot}$ and overdamped region $\gamma\gg\Omega_{\rm rot}$ respectively.
@@
Fig. 8 shows the dependence of the dephasing factor $\Gamma$ concerning different damping rates $\gamma$ and torsion noise amplitude $A$, where the parameters are chosen as $t_{a}=0.25$ s and $t_{e}=0$ s and $\Omega_{\rm rot}=2\pi\times 1$ Hz. As is shown, $\Gamma$ increases as $A$ increases or $\gamma$ decreases. In an experiment, if one requires an upper bound of $\Gamma=0.01$, then $A$ and $\gamma$ should be chosen in the region on the left side of the cyan region in the figure. In particular, $A$ should be roughly smaller than $10^{-10}\,\rm Hz^{3}$ and $10^{-6}\,\rm Hz^{3}$ for $\gamma=10^{-10}$ Hz and $10^{-2}$ Hz.
Fig. 9 shows the upper bound of the torsion noise amplitude $A$ to obtain a dephasing factor $\Gamma$ smaller than 0.01, where the parameters are also chosen as $t_{a}=0.25$ s and $t_{e}=0$ s. As is shown, the restriction on the upper bound of $A$ is very stringent near the resonance region $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}=2\pi\times 1$ Hz and its harmonic resonance $\Omega_{\rm rot}\sim 2n\pi/T_{\rm tot}$ (with $n$ a positive integer). In particular, $A$ has to be smaller than $10^{-10}\,\rm Hz^{3}$ for $\gamma=10^{-10}$ Hz, and it has to be smaller than $10^{-8}\,\rm Hz^{3}$ for $\gamma=10^{-2}$ Hz. On the other hand, the value of $A$ is less severely constrained outside the resonance region. For a damping rate larger than $10^{-4}$ Hz, $A$ can be larger than $10^{-5}\,\rm Hz^{3}$.
In a recent simulation work [52], a relationship between the torsion noise amplitude $A$ and the superposition size $\Delta x$ has been obtained as $\Delta xA\sim 10^{-11}\,\rm m\cdot Hz^{3}$. Since the superposition size discussed in this paper is $\Delta x=a_{m}t_{a}^{2}=11.2\,\rm\mu m$, then the noise amplitude is $A\sim 10^{-6}\,\rm Hz^{3}$, which is smaller than the restricted bound $A_{\rm bound}\sim 10^{-5}\,\rm Hz^{3}$. In conclusion, as long as the damping rate is designed to be larger than $10^{-4}$ Hz and the intrinsic frequency $\Omega_{\rm rot}$ is designed outside the resonance region, the dephasing factor $\Gamma$ of the interferometer will be smaller than $0.01$.
Figure 8: The dephasing factor $\Gamma$ with respect to different damping rates $\gamma$ and torsion noise amplitudes $A$. The transfer function is given by Eq. (13) with parameters $t_{a}=0.25$ s and $t_{e}=0$ s, and the intrinsic frequency is chosen as $\Omega_{\rm rot}=2\pi\times 1$ Hz. If $\Gamma$ is required to be smaller than 0.01, then the parameters should be chosen on the left side of the cyan region in the figure.
@@
VII Summary
In this paper, we investigated inertial torsion noise (ITN) in the context of matter-wave interferometry. In section II, we discussed the physical interpretation of the ITN as stochastic torsion of the suspended experimental apparatus. The consequence of such noise is a random differential phase $\delta\phi$ on the paths of the interferometer resulting in dephasing. The coherences, i.e., the off-diagonal terms of the density matrix, decay as $\mathbb{E}[\mathrm{e}^{\mathrm{i}\delta\phi}]=\mathrm{e}^{-\Gamma}$, where the decay factor is precisely the variance $\Gamma=\mathbb{E}[(\delta\phi)^{2}]$. In other words, ITN induces a loss of interferometric visibility.
The Lagrangian of the ITN is derived in Appendix A in the language of general relativity. We first constructed the rotating metric starting from Fermi normal coordinates, then computed the Lagrangian of a test mass under this metric in a non-relativistic limit. There are two terms in the Lagrangian, one of which is the centrifugal force, while the other corresponds to the Coriolis and Euler forces. Assuming the small angle approximation, and placing the superposition along the x-axis, the Lagrangian of the ITN in Eq.(5) can then be derived.

### v2

@@ IV Power Spectrum Density of Inertial Torsion Noise
There is also a peak of the PSD of ITN, of which the resonance frequency translates from $\Omega_{\rm rot}$ to $2\Omega_{\rm rot}$, which is a common property for a squared noise  [^3] This property can be understood as follows. Since $S_{\dot{\Theta}\dot{\Theta}}$ has a peak at $\Omega_{\rm rot}$, then it can be written as a $\delta$-function plus some small function $o(\omega)$: $S_{\dot{\Theta}\dot{\Theta}}(\omega)=S_{0}\delta(\omega-\Omega_{\rm rot})+o(\omega),$ where $S_{0}$ is the amplitude of the peak. Then the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}$ gives $\begin{split}S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)&=S_{0}^{2}\int\delta(\omega^{\prime}-\Omega_{\rm rot})\delta(\omega-\omega^{\prime}-\Omega_{\rm rot})\mathrm{d}\omega^{\prime}+o^{\prime}(\omega)\\ &=S_{0}^{2}\delta(\omega-2\Omega_{\rm rot})+o^{\prime}(\omega),\end{split}$ which means the peak position locates at $2\Omega_{\rm rot}$. In Appendix B, we offer another method to understand this property in the time domain.. In a small damping limit $\gamma_{\rm rot}\ll\Omega_{\rm rot}$, the peak value of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(2\Omega_{\rm rot})\approx\frac{A^{2}}{2\gamma^{3}}.$ (21)
It is notatble that
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(2\Omega_{\rm rot}\pm\gamma)\approx\frac{A^{2}}{4\gamma^{3}}.$ (22)
So the frequency at which the PSD equals to the half peak value are $\omega=2\Omega_{\rm rot}\pm\gamma$ and the corresponding full width at half maximum (FWHM) is $2\gamma$. The radio between the bandwidth and the peak value is known as the quality factor (Q-factor) of the PSD, which characterizes the sharpness of the peak. Then the Q-factor of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is
$Q_{\rm ITN}=\frac{2\Omega_{\rm rot}}{2\gamma}=\frac{\Omega_{\rm rot}}{\gamma},$ (23)
which is exactly the same as the Q-factor of $S_{\Theta\Theta}(\omega)$.
@@ V Dephasing Factor and Experiment Parameters Constrain
Based on the result of the transfer function in Eq. (11) or Eq. (14), and the power spectrum density (20) of the inertial torsion noise, the dephasing factor $\Gamma$ can be computed through the integral in Eq. (8). Note that the analytic solution to the integral is very complicated, so we compute it numerically. It is noteworthy that a resolution of $\omega_{\rm min}=2\pi/T_{\rm tot}$ exists as a cutoff in frequency space for the numerical calculation. Physically, this cutoff indicates that a low-frequency signal or noise will not be measurable by the experiment if the total experiment time is shorter than a single period of such signal or noise.
Figure 5: Dephasing factor $\Gamma$ as a function of the intrinsic torsion frequency $\Omega_{\rm rot}$. The parameters are chosen as $m_{0}=10^{-15}$ kg, $\eta=10^{4}$ T/m, $t_{a}=0.25$ s and $t_{e}=0$ s for the interferometer, and $\gamma=10^{-10}$ Hz, $I=10\,\rm kg\cdot m^{2}$ and $A=10^{-10}\,\rm Hz^{4}/Hz$ for the PSD of ITN. There is a resonance between the ITN and the transfer function with the first dominant peak at $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}\sim 2\pi\times 1$ Hz, and additional smaller peaks visible in the range upto 10. In the low frequency limit and high frequency limit, the dephasing factor $\Gamma$ is approximately independent of the intrinsic frequency $\Omega_{\rm rot}$.
@@
Fig. 5 shows the dependence of $\Gamma$ on the intrinsic torsion frequency $\Omega_{\rm rot}$ of the experiment apparatus, where the parameters of $F(\omega)$ are chosen as $m_{0}=10^{-15}$ kg, $t_{a}=0.25$ s and $t_{e}=0$ s, and the parameters of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ are $A=10^{-10}\,\rm Hz^{4}/Hz$ and $\gamma=10^{-10}$ Hz. As is shown, there is a resonance between the ITN and the transfer function near $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}=2\pi\times 1$ Hz and some harmonic resonances. Note that near the resonance peak, the precision of the numerical calculation is limited because the value is highly related to the sample points in frequency space.
@@
When $\Omega_{\rm rot}$ is much larger than $2\pi\times 10^{0}$ Hz, $\Gamma$ becomes approximately independent on $\Omega_{\rm rot}$. On the other hand, when $\Omega_{\rm rot}$ is much smaller than $2\pi\times 10^{-1}$ Hz, the peak will be below the frequency cut-off $\omega_{\rm min}$, so the dephasing factor will be tiny. For an actual experiment, the low intrinsic frequency limit is expected, so the parameters $d$, $l$, $\mathbb{G}$ and $I$ in (18) have to be carefully designed to ensure $\Omega_{\rm rot}$ much smaller than the resonance frequency $2\pi/T$ of the interferometer.
@@
Fig. 6 shows the dependence of $\Gamma$ on the damping rate $\gamma$. As is shown, when the damping rate is very low, i.e. $\gamma\ll\Omega_{\rm rot}$, the dephasing parameter $\Gamma$ is approximately proportional to $\gamma^{-1}$ due to the $\gamma^{-1}$ factor in (20).
However, when $\gamma$ increases after a critical point $\gamma\sim\Omega_{\rm rot}$, the dephasing $\Gamma$ is approximately $\gamma^{-3}$. This is because the PSD of ITN at the high damping rate limit $\gamma\gg\Omega_{\rm rot}$ is approximately
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)\approx A^{2}\frac{4\omega^{4}+4\gamma_{\rm rot}^{2}\omega^{2}}{\gamma(\omega^{2}+\gamma^{2})(\omega^{4}+4\gamma^{2}\omega^{2})}\sim\frac{A^{2}}{\gamma^{3}}.$ (24)
Physically it can be interpreted as an overdamped oscillator. In particular, the apparatus’s dynamical equation in Eq. (17) describes a damped oscillator under a randomly driven force. When the damping rate is larger than a critical value $\gamma>\Omega_{\rm rot}/2$, the system decays with no oscillation, known as overdamped. In the overdamped region, as the damping rate increases, the system decays to the equilibrium faster, so the random force term affects the system less. Finally, the dephasing parameter $\Gamma$ decreases as the damping rate $\gamma$ increases.
Figure 7: The dephasing parameter $\Gamma$ with respect to different damping rates $\gamma$ and torsion noise amplitudes $A$. The transfer function is given by Eq. (11) with parameters $t_{a}=0.25$ s and $t_{e}=0$ s, and the intrinsic frequency is chosen as $\Omega_{\rm rot}=2\pi\times 1$ Hz. $\Gamma$ is required smaller than some thresholds in different situations, then the parameters should be chosen on the left side of the critical lines in the figure, where the dotted and dashed lines represent $\Gamma=10^{-6}$ and $\Gamma=0.01$ for gravimeters and QGEM experiment in respect.
Fig. 7 shows the dependence of the dephasing parameter $\Gamma$ concerning different damping rates $\gamma$ and torsion noise amplitude $A$, where the parameters are chosen as $t_{a}=0.25$ s and $t_{e}=0$ s and $\Omega_{\rm rot}=2\pi\times 1$ Hz. As is shown, $\Gamma$ increases as $A$ increases or $\gamma$ decreases. As is discussed in section 2, the gravimeter and QGEM experiments require an upper bound of $\Gamma$ as $10^{-6}$ and $0.01$, then $A$ and $\gamma$ should be chosen in the region on the left side of the dotted and dashed critical line respectively in the figure.
@@
Fig. 8 shows the upper bound of the torsion noise amplitude $A$ to obtain a dephasing parameter $\Gamma$ smaller than 0.01, where the parameters are also chosen as $t_{a}=0.25$ s and $t_{e}=0$ s. Since $\Gamma$ is proportional to $A^{2}$, so one can also use Fig. 8 to analyze other threshold of $\Gamma$ by multiplying a factor on $A$. For example, one can analyze the case $\Gamma<10^{-6}$ for gravimeters by multiplying $10^{-2}$ to the upper bound of $A$.
As is shown, the restriction on the upper bound of $A$ is very stringent near the resonance region $\Omega_{\rm rot}\sim 2\pi/T_{\rm tot}=2\pi\times 1$ Hz and its harmonic resonance $\Omega_{\rm rot}\sim 2n\pi/T_{\rm tot}$ (with $n$ a positive integer). In particular, $A$ has to be smaller than $10^{-10}\,\rm Hz^{3}$ for $\gamma=10^{-10}$ Hz, and it has to be smaller than $10^{-8}\,\rm Hz^{3}$ for $\gamma=10^{-2}$ Hz. On the other hand, the value of $A$ is less severely constrained outside the resonance region. For a damping rate larger than $10^{-4}$ Hz, $A$ can be larger than $10^{-5}\,\rm Hz^{3}$.
In a recent simulation work [56], a relationship between the torsion noise amplitude $A$ and the superposition size $\Delta x$ has been obtained as $\Delta xA\sim 10^{-11}\,\rm m\cdot Hz^{3}$. Since the superposition size discussed in this paper is $\Delta x=a_{m}t_{a}^{2}=11.2\,\rm\mu m$, then the noise amplitude is $A\sim 10^{-6}\,\rm Hz^{3}$, which is smaller than the restricted bound $A_{\rm bound}\sim 10^{-5}\,\rm Hz^{3}$. In conclusion, as long as the damping rate is designed to be larger than $10^{-4}$ Hz and the intrinsic frequency $\Omega_{\rm rot}$ is designed outside the resonance region, the dephasing parameter $\Gamma$ of the interferometer will be smaller than $0.01$.
VI Summary
In this paper, we investigated inertial torsion noise (ITN) in the context of matter-wave interferometry, and some of our highlights and conclusions are summarized below.
In section II, we explained the physical interpretation of the ITN and briefly reviewed the dephasing effect of generic noises on matter-wave interferometers. The key point is that the ensemble average of a random phase implies a decay factor $\mathbb{E}[\mathrm{e}^{\mathrm{i}\delta\phi}]=\mathrm{e}^{-\Gamma}$ on the off-diagonal terms of the density matrix, where the dephasing parameter is exactly the variance of the noise $\Gamma=\mathbb{E}[(\delta\phi)^{2}]$.